OpenStudy (anonymous):

What is the nuclear binding energy for uranium-238 in joules? Assume the following: Mass defect = 3.2008 x 10-27 kilogramsUse E = mc2, with c = (3 x 108 m/s)2 A. 0.28807 x 10-12 joules B. 2.8807 x 10-12 joules C. 2.8807 x 10-10 joules D. 2.8807 x 10-8 joules E. 3.2008 x 1027 joules

OpenStudy (souvik):

use calculator and find Energy.. the formula is given in the question..........E=mc^2 here m=mass defect and c=the speed of light in vacuum

OpenStudy (anonymous):

do i plug this into the equation E = mc^2=7.56 x 10 to the power of 13

OpenStudy (souvik):

3.2008*10^-27*(3*10^8)^2=?